Three times the smallest of three consecutive even integers is two more than twice the largest. Find the integers.
Answer by addingup(3677)

3(x) = 2(x+4)+2
3x = 2x+8+2
3x = 2x+10
x = 10
So your even integers are 10, 10+2, 10+4 = 10, 12, 14
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
The problem says:
3* the smallest is 2+ than 2* the largest
3(10) = 2+(2*14) Use your calculator to see that we have the correct answer. Try doing it in your head first.
J